# Physics-Informed Multi-Agent Reinforcement Learning for Distributed Multi-Robot Problems

### v1

@@ V-B The environment
The second modification over the soft actor-critic algorithm is in the collection of experience from the environment. The environment is determined by the desired task to be solved, with examples found in Fig. 1. The main difference with other multi-agent reinforcement learning works is that, to build the replay buffer that stores the trial and error experience used for training the control policies, we take into account all the interactions among the robots. Differently from centralized-training decentralized-deployment approaches, where the experience of each agent is recorded independently of the other agents, we record for each experience all the robot states, actions, and interaction graph together. Thanks to that, we keep track of the correlation among robots during training. This also allows us to condition the trained control policies on the available information at each robot, therefore only providing the information robots will have access during deployment.
@@ V-C The reward
The reward function can be shared across robots and include global terms because the actor is the whole multi-robot network. This is important because, from the perspective of the soft actor-critic algorithm during training, the whole robot network is a single centralized agent. Other state-of-the-art multi-agent reinforcement learning approaches (see Sec. II) require a factorization of the reward function because each agent is considered as an isolated learning unit which does not exploit neighboring information during the execution of the control policy. In contrast, in our approach the shared distributed control policy is simultaneously learned at all the robots used in training, and since the experience record the underlying communication graph, we can seamlessly associate global rewards with distributed cooperative control policies through communication. It is worth to note that, as in any classical soft actor-critic approach, the maximization objective is changed to include an entropy term ${\cal H}(\bm{\Pi}_{{\bm{\theta}}}(\mathbf{a}|\mathbf{s}))$ that measures the entropy of the control policy. This entropy term, weighted by a temperature parameter $\alpha>0$, trades off exploration (high value) and exploitation (low value). The value of $\alpha$ changes over time according to a gradient descend law which manages automatically the exploration/exploitation dichotomy. The details of this automatic temperature adjustment rule can be found in [73].
@@ V-D The critic
Finally, regarding the critic, its main purpose is to learn the action-value function of the environment, encoded in $Q_{\bm{\Pi}_{\bm{\theta}}}(\mathbf{s},\mathbf{a})$. This approximation steers the training of the control policy towards the maximization objective. Nevertheless and importantly, the critic is only used during training. Thus, it is not necessary to design it to be distributed and an existing centralized neural network architecture can be used. In this work, we opt for a multi-layer perceptron with parameters $\bm{\vartheta}$ to learn $Q^{\bm{\vartheta}}_{\bm{\Pi}_{\bm{\theta}}}(\mathbf{s},\mathbf{a})$. As with the reward, since, from the perspective of the soft actor-critic algorithm, the single agent is the whole robot network described by the port-Hamiltonian dynamics, a single centralized critic can learn the appropriate action-value function to condition the training of the distributed control policy.
@@ VI Results
To assess our physics-informed multi-agent reinforcement learning approach, we present six multi-robot scenarios. The first three are extracted from the VMAS simulator [74] and the last three are adaptations from the MPE simulator [76, 57] that can be found in [60]:
@@
The first three scenarios are used for ablation studies, while the other three are used to compare our method with other state-of-the-art approaches.
All robots have a communication radius $r_{comm}>0$ which allows to exchange information with $1$-hop neighbors. The communication radius and other hyperparameters of the soft actor-critic algorithm are specified in Appendix B. Each robot observes its position, velocity, position and velocity of the landmarks or objects of interest (e.g., the box in reverse transport). The particularities of the observation space of each robot can be found in [74, 57, 60]. In some scenarios, the robot observation vector changes its dimension depending on the number of robots. Specifically, in the food collection, grassland and adversarial scenarios the number of landmarks changes with the number of robots. Therefore, to accommodate the proposed self-attention port-Hamiltonian neural network with an observation vector which may change its size, we use an additional neural network to pre-process the observation vector. In particular, we concatenate a self-attention layer with a dense layer which receives, as input, the food landmarks’ positions and outputs a feature vector of constant dimension that is used as part of the state vector to build $\mathbf{S}_{t}^{i}$. The details can be found in Appendix A. Supplementary material can be found in our repository111 https://github.com/EduardoSebastianRodriguez/phMARL.
@@
Figure 5: Examples of multi-robot scenarios for different initial conditions and number of robots. It is interesting to see that some aspects of the environment do not scale with the team size, e.g., the size and weight of the box in the reverse transport scenario, the number of hot spots in the sampling scenario or the size of the arena in the navigation scenario.
@@ VI-A Ablation results
We conduct ablation studies using the reverse transport, sampling and navigation scenarios. We compare the proposed self-attention port-Hamiltonian neural network (pH-MARL) with a classical multi-layer perceptron (MLP), a modular self-attention-based neural network (MSA) and a graph- and attention-based neural network (GSA). The implementation details can be found in Appendix A.
@@ VI-B Comparative results
We compare the performance of our proposed physics-informed multi-agent reinforcement learning approach with other state-of-the-art multi-agent reinforcement learning approaches. In particular, we compare with: Multi-Agent Deep Deterministic Policy Gradient (MADDPG) [57], Mean Field Actor Critic (MFAC) [59], Evolutionary Population Curriculum (EPC) [77] and Distributed multi-Agent Reinforcement Learning with One-hop Neighbors (DARL1N) [60]. The comparison results reported in this paper are extracted from [60].
@@
Next, we assess the scalability of our proposed approach in this second series of scenarios. We train the other state-of-the-art approaches for each specific number of robots, ranging from $3$ to $48$ robots. In contrast, we train pH-MARL $n=4$ robots and evaluate the trained control policies with the other number of robots, without further training. Fig. 6 demonstrates that pH-MARL achieves better or similar results than the other approaches without further training the control policies. When the team size is close to the one used during training ($n=3,6,12$) pH-MARL surpasses the other methods in the three scenarios, proving the accuracy of combining a physics-informed description of the multi-robot team with self-attention mechanisms. When the team size is much greater than the one used during training ($24,48$), pH-MARL outperforms (MADDPG, MFAC) or achieves similar results (EPC, DARL1N) as the state-of-the-art. Notably, pH-MARL outperforms all the state-of-the-art methods in the adversarial scenario even when $n=48$, which is significant considering that they are all trained with $n=48$ robots except pH-MARL, which is trained with $n=4$ robots.
@@
Figure 6: Comparison of our proposed physics-informed multi-agent reinforcement learning approach with other state-of-the-art approaches. pH-MARL is only trained with $n=4$ robots and deployed with different team sizes, while the state-of-the-art control policies are trained for each specific number of robots.

### v3

@@ V-B The environment
The second modification over the soft actor-critic algorithm is in the collection of experiences from the environment. The environment is determined by the desired task to be solved, with examples found in Fig. 1. The main difference with other multi-agent reinforcement learning works is that, to build the replay buffer that stores the trial and error experience used for training the control policies, we take into account all the interactions among the robots. Differently from centralized-training decentralized-deployment approaches, where the experience of each agent is recorded independently of the other agents, we record for each experience all the robot states, actions, and interaction graph together. Thanks to that, we keep track of the correlation among robots during training. This also allows us to condition the trained control policies on the available information at each robot, therefore only providing the information robots will have access during deployment.
@@ V-C The reward
The reward function can be shared across robots and include global terms because the actor is the whole multi-robot network. This is important because, from the perspective of the soft actor-critic algorithm during training, the whole robot network is a single centralized agent. Other state-of-the-art multi-agent reinforcement learning approaches (see Sec. II) require a factorization of the reward function because each agent is considered as an isolated learning unit which does not exploit neighboring information during the execution of the control policy. In contrast, in our approach the shared distributed control policy is simultaneously learned at all the robots used in training, and since the experience records the underlying communication graph, we can seamlessly associate global rewards with distributed cooperative control policies through communication. It is worth to note that, as in any classical soft actor-critic approach, the maximization objective is changed to include an entropy term ${\cal H}(\bm{\Pi}_{{\bm{\theta}}}(\mathbf{a}|\mathbf{s}))$ that measures the entropy of the control policy. This entropy term, weighted by a temperature parameter $\alpha>0$, trades off exploration (high value) and exploitation (low value). The value of $\alpha$ changes over time according to a gradient descend law which manages automatically the exploration/exploitation dichotomy. The details of this automatic temperature adjustment rule can be found in [80].
@@ V-D The critic
Finally, regarding the critic, its main purpose is to learn the action-value function of the environment, encoded in $Q_{\bm{\Pi}_{\bm{\theta}}}(\mathbf{s},\mathbf{a})$. This approximation steers the training of the control policy towards the maximization objective. Nevertheless and importantly, the critic is only used during training. Thus, it is not necessary to design it to be distributed and an existing centralized neural network architecture can be used. In this work, we opt for a multi-layer perceptron with parameters $\bm{\vartheta}$ to learn $Q_{\bm{\vartheta},\bm{\Pi}_{\bm{\theta}}}(\mathbf{s},\mathbf{a})$. As with the reward, since, from the perspective of the soft actor-critic algorithm, the single agent is the whole robot network described by the port-Hamiltonian dynamics, a single centralized critic can learn the appropriate action-value function to condition the training of the distributed control policy.
@@ VI Results
To assess our physics-informed multi-agent reinforcement learning approach, we present seven multi-robot scenarios. The first three are extracted from the VMAS simulator [81]. The next three are adaptations from the MPE simulator [83, 59] that can be found in [65]. The last scenario comes from the Multi-Agent MuJoCo benchmark [2]. The overview of the scenarios are detailed in the following:
@@
  7. g.
    6x1-Half Cheetah: a robotic cheetah with two legs and 6 joints, where each joint is a different agent. The multi-agent team is distributed in the sense that each agent can communicate only with the adjacent joints, enforced by a ring graph topology. The task is to make the robot run as fast as possible. The reward is global and is composed by a first term that favours forward movements and a second term that penalizes too large actions.
The first three scenarios are used for ablation studies, where the main goal is to evaluate how our physics-informed policy parameterization improves upon existing standard policy parameterizations. The second three are used to compare our method with other state-of-the-art multi-agent reinforcement learning approaches that are not physics-informed. The last scenario is used to validate our approach in a realistic robotic platform.
All robots have a communication radius $r_{comm}>0$ which allows to exchange information with $1$-hop neighbors. The communication radius and other hyperparameters of the soft actor-critic algorithm are specified in Appendix B. Each robot observes its position, velocity, position and velocity of the landmarks or objects of interest (e.g., the box in reverse transport). The particularities of the observation space of each robot can be found in [81, 59, 65]. In some scenarios, the robot observation vector changes its dimension depending on the number of robots. Specifically, in the food collection, grassland and adversarial scenarios the number of landmarks changes with the number of robots. Therefore, to accommodate the proposed self-attention port-Hamiltonian neural network with an observation vector which may change its size, we use an additional neural network to pre-process the observation vector. In particular, we concatenate a self-attention layer with a dense layer which receives, as input, the food landmarks’ positions and outputs a feature vector of constant dimension that is used as part of the state vector to build $\mathbf{S}_{t}^{i}$. The details can be found in Appendix A. Supplementary material can be found in our repository111 https://github.com/EduardoSebastianRodriguez/phMARL and the supplementary video [^2] https://youtu.be/pSzP3LBVyZg.
@@
Figure 5: Examples of multi-robot scenarios for different initial conditions and numbers of robots. In the reverse transport scenario, the red box turns green when it reaches the desired goal (green dot), and the robots are depicted as small blue dots. In the sampling scenario, LiDAR measurements are depicted as points forming a star pattern centered at the position of the robot, and regions already explored are shown in orange. In the navigation scenario, evolution over time of the trajectories is colored from pink ($t=0$) to blue ($t=t_{f}$), whereas the goals are depicted as black dots. It is interesting to see that some aspects of the environment do not scale with the team size, e.g., the size and weight of the box in the reverse transport scenario, the number of hot spots in the sampling scenario or the size of the arena in the navigation scenario.
@@ VI-A Ablation results
We conduct ablation studies using the reverse transport, sampling and navigation scenarios. The goal is to evaluate whether our physics-informed policy parameterization improves upon existing standard policy parameterizations. We compare the proposed self-attention port-Hamiltonian neural network (pH-MARL) with a classical multi-layer perceptron (MLP), a modular self-attention-based neural network (MSA) and a graph- and attention-based neural network (GSA). The implementation details can be found in Appendix A.
@@ VI-B Comparative results
We compare the performance of our proposed physics-informed multi-agent reinforcement learning approach with other state-of-the-art multi-agent reinforcement learning approaches that are not physics-informed. In particular, we compare with: Multi-Agent Deep Deterministic Policy Gradient (MADDPG) [59], Mean Field Actor Critic (MFAC) [64], Evolutionary Population Curriculum (EPC) [84], Distributed multi-Agent Reinforcement Learning with One-hop Neighbors (DARL1N) [65] and Multi-Agent Proximal Policy Optimization (MAPPO) [60]. The comparison results of the first four methods reported in this paper are extracted from [65].
@@
Next, we assess the scalability of our proposed approach in this second series of scenarios. We train the other state-of-the-art approaches for each specific number of robots, ranging from $3$ to $48$ robots. In contrast, we train pH-MARL $n=4$ robots and evaluate the trained control policies with the other number of robots, without further training. Fig. 6 demonstrates that pH-MARL achieves better or similar results than the other approaches without further training the control policies. When the team size is close to the one used during training ($n=3,6,12$) pH-MARL surpasses the other methods in the three scenarios, proving the accuracy of combining a physics-informed description of the multi-robot team with self-attention mechanisms. In this sense, MAPPO in the grassland scenario with $n=6$ robots is the only case that obtains better results than pH-MARL. When the team size is much greater than the one used during training ($24,48$), pH-MARL outperforms (MADDPG, MFAC) or achieves similar results (EPC, DARL1N, MAPPO) as the state-of-the-art. Notably, pH-MARL outperforms all the state-of-the-art methods in the adversarial scenario even when $n=48$, which is significant considering that they are all trained with $n=48$ robots except pH-MARL, which is trained with $n=4$ robots. Regarding MAPPO, despite its strong performance in some cases (e.g., food collection with $n=3$ robots, grassland with $n=6$ robots or grassland with $n=48$ robots), it generally exhibits poor results when the number of robots is larger than $10$ robots. Besides, MAPPO is unable to learn successful policies that avoid collisions and find the food resources for any of the teams in the adversarial scenario.
@@
Figure 6: Comparison of our proposed physics-informed multi-agent reinforcement learning approach with other state-of-the-art approaches. To measure performance, we use the averaged cumulative reward ¯$\bm{\bar{}}{R}=\frac{1}{n}R=\frac{1}{n}\sum_{\tau=0}^{\tau_{\max}-1}r(\mathbf{s}_{\tau},\mathbf{a}_{\tau})$. pH-MARL is only trained with $n=4$ robots and deployed with different team sizes, while the state-of-the-art control policies are trained for each specific number of robots. For team sizes similar to the one used in training, our policy outperforms the other approaches. For team sizes much larger than the training team size, our policy still outperforms or achieves similar results as the state-of-the-art approaches.
Regarding the computational cost, it is worth noting that, given the same scenario, optimizer and replay buffer hyperparameters, the main differences among methods come from the parameterization of the actor. In this sense, our approach involves greater training times per sample and robot compared to other non-physics-informed methods. Specifically, the increase in computational burden comes from the differentiation of the Hamiltonian $\partial H_{{\bm{\theta}}}(\mathbf{x})/\partial\mathbf{x}_{j}$ at each time step. However, by restricting the space of admissible policies, we have better sample efficiency during training compared to non-physics-informed parameterizations. Therefore, the increase of computational burden per sample and robot is compensated by the sample efficiency of the method, leading to similar overall computational times in robot tasks during training. Besides, as it has been observed in the scalability results, our approach needs a very small number of robots to reproduce or even outperform the other methods when they are trained with larger numbers of robots, saving additional computation. Importantly, the computational cost during deployment scales with the number of neighbors rather than the network size, therefore preserving the desirable properties coming from a distributed multi-robot setting. Alternatively, to improve computational efficiency, one could consider a neural network estimator that learns to infer $\partial H_{{\bm{\theta}}}(\mathbf{x})/\partial\mathbf{x}_{j}$ from $\mathbf{x}^{i}_{{\cal N}^{i,k}}$, which bypasses automatic differentiation of $H_{{\bm{\theta}}}(\mathbf{x})$ and the three message-passing communication protocol. Avoiding automatic differentiation to reduce inference time is an active area of research [85], and we leave its exploration to optimize the training time of our method for future work.
(a) Reverse transport (b) Sampling (c) Navigation
Refer to caption Refer to caption Refer to caption
Figure 7: Sensitivity analysis of pH-MARL with respect to the number of robots at training. To measure performance, we use the averaged cumulative reward ¯$\bm{\bar{}}{R}=\frac{1}{n}R=\frac{1}{n}\sum_{\tau=0}^{\tau_{\max}-1}r(\mathbf{s}_{\tau},\mathbf{a}_{\tau})$. Compared to the nominal value ($n=4$), increasing the number of training robots only leads to greater sample complexity, whereas decreasing the number of training robots does not allow to learn sufficiently varied behaviors to scale with the number of robots.
VI-C Sensitivity analysis
Regarding the evaluation of the training scalability of our approach, we have conducted a series of experiments using the same scenarios of Section VI-A and Section VI-B. Given those scenarios, we have trained our ph-MARL with an increasing number of robots, namely, $n=\{2,4,8,12\}$ (see Fig. 7). The first interesting conclusion is that the increase in number of robots does not necessarily lead to better performance because the interactions among robots always happen at the robot neighborhood level. In the reverse transport scenario, more robots inside the cage implies a greater number of robot contacts with the cage and, therefore, a drop in averaged cumulative reward. In the sampling scenario, a greater number of robots leads to an improvement in performance because, for the same time horizon, robots can explore the arena faster and localize the informative regions faster. In the case of the navigation scenario, due to space constraints, an increase in the number of robots during training leads to a significant drop in averaged cumulative reward because the space available to each robot is smaller, so more collisions happen (specially with $n=12$). In all the scenarios, $n=2$ is a too small multi-robot team size to capture the richness of distributed multi-robot interactions, so the training process is unable to learn a successful policy in terms of averaged cumulative reward.
Therefore, in general, the optimal number of training robots depends on constraints of the scenario and the task. Nonetheless, as a rule of thumb, increasing the number of training robots towards a large multi-robot team only leads to greater sample complexity, whereas decreasing the number of training robots does not allow to learn sufficiently varied behaviors to scale with the number of robots. Henceforth, we conclude that our approach is best suited for training with a small number of training robots (relative to the task/scenario) but still sufficiently large to capture all possible neighborhood interactions.
From a behavioral analysis perspective, the results in Fig. 7 suggest that the learned distributed multi-robot policies benefit from small neighborhoods, irrespective of the network size. In scenarios with a fixed arena size, an increase in the number of robots indirectly leads to an increase in neighborhood sizes, with a subsequent drop in averaged cumulative reward both in training and evaluation (see Fig. 6). The appropriate number of training robots depends on the scenario, e.g., in the reverse transport scenario the best performance is achieved with $n=4$ while in the sampling scenario this is obtained with $n=8$. In any case, as concluded above, our approach requires a sufficiently large team size to capture a variety of local interactions. On the other hand, different from other centralized-training decentralized-execution approaches, our approach considers a single set of policy parameters that is invariant to the number of robots in the team due to its distributed structure. Therefore, behavioral analysis to understand individual influence in the overall team reduces to understanding how different neighborhood sizes affect overall performance, which is already given by the sensitivity analysis with respect to the number of training robots for a fixed environment size, and the scalability tests shown in Fig. 4 and Fig. 6.
VI-D Validation on a realistic robot platform
TABLE II: Physical properties of the joints of the MuJoCo Half Cheetah.
Joint Type Range [rad] Damping [Ns/m] Stiffness [N/m]
Back thigh hinge $[-0.52,1.05]$ $6$ $240$
Back shin hinge $[-0.785,0.785]$ $4.5$ $180$
Back foot hinge $[-0.4,0.785]$ $3$ $120$
Front thigh hinge $[-1,0.7]$ $4.5$ $180$
Front shin hinge $[-1.2,0.87]$ $3$ $120$
Front foot hinge $[-0.5,0.5]$ $1.5$ $60$
We further validate pH-MARL in a realistic robot setting using MuJoCo [86]. MuJoCo is a general purpose physics engine that replicates multi-joint dynamics with contact, allowing for fast yet realistic physical behaviors. Specifically, we use the Multi-agent MuJoCo benchmark [2], where the task is to learn coordination policies for multi-joint robots that make them walk. Our approach was evaluated using the 6x1-Half Cheetah environment (see Fig. 1e). Since our fundamental contribution is the physics-informed distributed policy parameterization, we integrated our policy design in a Trust-Region Policy Optimization (TRPO) reinforcement learning algorithm [87]. This demonstrates the flexibility of the physics-informed policy parameterization in the sense that it can be integrated in different reinforcement learning algorithms. The benchmark provides results from state-of-the-art multi-agent reinforcement learning approaches for comparison, including Heterogeneous Agent Trust Region Policy Optimization (HATRPO) and Heterogeneous-Agent Proximal Policy Optimization (HAPPO) [62], MAPPO, IPPO [66], and MADDPG. Importantly, in all these state-of-the-art methods, the policy of each agent/joint depends on the global observation vector, whereas our policy only depends on the neighboring information, enforcing a ring topology.
Figure 8: Average episode reward obtained by ph-MARL and other MARL approaches in the 6x1-Half Cheetah environment.
The 6x1-Half Cheetah model is composed of $n=6$ joints, corresponding to: {back thigh, back shin, back foot, front thigh, front shin, front foot}. The input to each joint is a torque with [Nm] units. The local observation of each agent is the angle, angular velocity, linear velocity, angular acceleration and linear acceleration of the joint; and the (x,z)-position, angular velocity and (x,z)-linear velocity of front tip. The physical properties of the joints follow those of the original MuJoCo Half Cheetah [86], detailed in Table II. Importantly, physical interactions like friction of the joints with the armature of the main body are computed numerically from the stiffness and damping of the joint. Stiffness is modeled as a spring with equilibrium at the nominal position of the joint, whereas the damping is modeled as a force linear in velocity.
Figure 9: Quantitative results of the pH-MARL policy for multi-robot navigation in the Robotarium simulator. From left to right, the following metrics are reported for different numbers of robots and communication quality (baseline means perfect communications): travel distance, travel time and success rate. All the metrics are the average over $10$ random runs.
(a) $n=4$ (b) $n=8$ (c) $n=12$ (d) $n=16$
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Qualitative results of the pH-MARL policy for multi-robot navigation in the Robotarium simulator. In (a), communication is perfect, whereas (b), (c) and (d) operate under imperfect communication. In the imperfect communication cases, magenta links denote delays, red links denote packet loss, yellow links denote disturbances, and green links denote unperturbed links.
As it can be observed in Fig. 8, our method, although less informed than the others due to the networked constraints, improves upon IPPO, MADDPG and MAPPO in average episode reward, and obtains a similar final average episode reward to HAPPO and HATRPO. Besides, by fully exploiting the physical dynamics of the robots, our method presents better sample efficiency than MADDPG, IPPO and MAPPO, and similar sample efficiency than HAPPO and HATRPO. In this sense, our method leverages the fact that the acceleration of the joints is available, so in the open-loop feed-forward term $\dot{\mathbf{x}}$ is known without explicit knowledge on the values of $\mathbf{J}(\mathbf{x})$, $\mathbf{R}(\mathbf{x})$ and $\partial{H}(\mathbf{x})/\partial\mathbf{x}$. This experiment shows that our approach is a promising method for multi-joint dynamical robots. The sample efficiency achieved by distributing the policy according to the imposed ring topology is reduced when the connectivity of the network increases. The ring topology restricts the neighborhood of joint $i$ to $\{i-1\mod 6,i,i+1\mod 6\}$, which implies a very sparse graph. This serves as a prior for how the other joints affect joint $i$ that restricts the class of policies that can be learned, leading to improved sample efficiency. In the extreme case where the topology is a fully connected graph, our approach would recover the centralized topology of the benchmarked methods. The performance in terms of average episode reward would, therefore, be at least the same, since more information is provided to the joints; however, the sample efficiency would be hampered by computing $6$ centralized policies, one per joint.
VI-E Real multi-robot experiments
Finally, we validate the sim-to-real transfer capabilities of pH-MARL using the Georgia Tech Robotarium [1]. The goal of this validation is twofold. First, in all previous sections we assumed perfect communication channels, whereas physical communication in real settings is subject to packet losses, disturbances in the messages and delays, leading to potentially adverse effects on the control policies. Second, we aim to evaluate how pH-MARL policies trained in an ideal simulated environment handle the transfer gap to real multi-robot settings subject to imperfect actuation in the differential-drive robot dynamics of the Robotarium robots. We pose a navigation problem with collision avoidance akin to the “Navigation” scenario used for ablations in Section VI-A. The only difference is that, to better represent the Robotarium arena, we train over a $3.2\times 2$m space and increase the physical radius of the agents to $11$cm to fit the size of the Robotarium robots. We train a pH-MARL navigation policy with $n=4$ as in Section VI-A in VMAS.
After that, we integrate the trained policy in the Robotarium simulator, designed to be as similar as possible to the Robotarium arena to guarantee that policies submitted to the real Robotarium platform comply with all requirements. Two settings are evaluated. The first one evaluates the pH-MARL policy with perfect communications with $r_{comm}=0.75$m. The second evaluates the pH-MARL policy with realistic communication, as follows.
  • Packet losses: An existing communication link has probability $0.1$ of being removed from the graph.
  • Communication disturbances: Messages sent through a communication channel have probability $0.1$ of being distorted by an additive zero-mean Gaussian noise with variance $0.05\mathbf{I}$.
  • Delays: Messages sent through a communication channel have probability $0.1$ of suffer a delay uniformly distributed in the range $[1,10]$ sampling times (sampling time is $0.033$s).
Fig. 9 compares the travel distance, travel time and success rate for the pH-MARL policy under perfect and imperfect communication for different number of robots, averaged over $10$ random runs each. The pH-MARL policy is able to scale to teams up to ×$\times 4$ larger than the one used for training, achieving constant travel distances and almost constant travel times with respect to $n$. The success rates are always above $95\%$, proving that the pH-MARL policy consistently transfers to different robot dynamics (training is done with a damped second-order integrator, whereas the Robotarium robots follow first-order differential-drive dynamics) and environment. Importantly, the pH-MARL policy is robust against imperfect communication, achieving the same results obtained with perfect communication. Qualitatively speaking, Fig. 10 shows that the robots achieve smooth trajectories and they reach their goals avoiding collisions.
Once we verified that the pH-MARL policy is scalable and robust against communication disturbances in the Robotarium simulator, we evaluated it in the real Robotarium arena under imperfect communication. All the parameters are the same, and every configuration of number of robots is averaged over $3$ random runs. As show in Fig. 11, the travel distance and travel time are consistent across number of robots, even though the space in the arena becomes increasingly tight with fleets of $12$ and $16$ robots. The case of $n=4$ is much easier than the others, and consequently the travel distance and time are smaller than in the other configurations. The success rate is always $100\%$, allowing us to conclude that our pH-MARL policy supports zero-shot sim-to-real transfer. This is reinforced by the qualitative results in Fig. 12, where robots follow smooth paths and avoid collisions akin to what the pH-MARL policy achieves in the Robotarium simulator.
Figure 11: Quantitative results of the pH-MARL policy for multi-robot navigation in the Robotarium arena under imperfect communication. From left to right: travel distance, travel time and success rate. All the metrics are the average over $3$ random runs.
(a) $n=4$ (b) $n=8$
Refer to caption Refer to caption
(c) $n=12$ (d) $n=16$
Refer to caption Refer to caption
Figure 12: Qualitative results of the pH-MARL policy for multi-robot navigation in the Robotarium arena under imperfect communication.